40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 400     Entry time: Fri Feb 3 16:50:02 2012
Author: Giordon Stark 
Type: DailyProgress 
Category: Coating Q 
Subject: New Setup (closer to what we planned out) 

 Lab set up update. Pictures and panorama attached.

Panorama (it's cool!): linky

We've taken apart the previous set up and included the vacuum chamber in it. This was done by slightly shifting/rotating the vacuum chamber in a way that the HV cables weren't super taut (or stretched) while allowing the beam to get through the whole thing. This part was tricky as the supports in the chamber block opposite halves of the windows such that, given a straight beam, the whole chamber has to be at an angle. The chamber also de-focused our beam so we added in a focusing lens at the far end of the set up before the polarizing beam splitter. Finally, we checked the output levels of the photodiodes and set up the wave plates so that we're at a balanced state (equal outputs out of both ends).

We should have the ESD by Monday, which means we'll have Alastair in the morning to help us quickly with the first set up so we're pretty confident in what to do.

Attachment 1: ExpSetup_Front.jpg  146 kB  | Hide | Hide all
ExpSetup_Front.jpg
Attachment 2: ExpSetup_Back.jpg  141 kB  | Hide | Hide all
ExpSetup_Back.jpg
ELOG V3.1.3-